Find sec(x) if sin(x)= -4/5 and 270 degrees < x < 360 degrees.

Can someone please help?

Answer by stanbon(75887)

Find sec(x) if sin(x)= -4/5 and 270 degrees < x < 360 degrees.
------------
sin(theta)= y/r
x=sqrt(r^2-y^2)
--------
Your problem:
y = -4
r=5
So, x=sqrt(25-16)=3
-----------
sec(theta) = r/x = 5/3
==========================
You can see all this if you draw an angle in the 4th quadrant;
let the distance from the origin be r=5; let the distance below
the x-axis be 4; then the 3rd side is x=3 and the sec(theta)=r/x=5/3
===================
Cheers,
Stan H.